OpenStudy (anonymous):

MEDAL AND FAN Rewrite f(x) = 2(x − 1)2 + 3 from vertex form to standard form. f(x) = 2x2 + 5 f(x) = 4x2 − 8x + 7 f(x) = 2x2 + 4x + 5 f(x) = 2x2 − 4x + 5

OpenStudy (jango_in_dtown):

\[2(x-1)^{2}+3=2(x^2-2x+1)+3=2x^2-4x+2+3\]

OpenStudy (jango_in_dtown):

=2x^2-4x+5

OpenStudy (jango_in_dtown):

@mariyakelly